# Integrating Multi-Physics Simulations and Machine Learning to Define the Spatter Mechanism and Process Window in Laser Powder Bed Fusion

### v1

@@ Abstract
Laser powder bed fusion (LPBF) has shown promise for wide range of applications due to its ability to fabricate freeform geometries and generate a controlled microstructure. However, components generated by LPBF still possess sub-optimal mechanical properties due to the defects that are created during laser-material interactions. In this work, we investigate mechanism of spatter formation, using a high-fidelity modelling tool that was built to simulate the multi-physics phenomena in LPBF. The modelling tool have the capability to capture the 3D resolution of the meltpool and the spatter behavior. To understand spatter behavior and formation, we reveal its properties at ejection and evaluate its variation from the meltpool, the source where it is formed. The dataset of the spatter and the meltpool collected consist of 50 % spatter and 50 % melt pool samples, with features that include position components, velocity components, velocity magnitude, temperature, density and pressure. The relationship between the spatter and the meltpool were evaluated via correlation analysis and machine learning (ML) algorithms for classification tasks. Upon screening different ML algorithms on the dataset, a high accuracy was observed for all the ML models, with ExtraTrees having the highest at 96 % and KNN having the lowest at 94 %. Furthermore, the ML model was tested on a different dataset generated using FLOW-3D, a commercial, less computationally expensive tool, but lacks the capability to simulate realistic spatter formation. Our ML model trained on OpenFOAM dataset was able to classify the region of the meltpool in the FLOW-3D datasets as spatter initiation location. Upon screening 283 FLOW-3D simulation experiments with varying power-velocity combinations, a process map based on spatter volume was established. The process map will serve as a guide towards understanding how the process parameters influence part properties, identifying an AM process window with minimal defects and ultimately enhance repeatability in part production.
@@ 1 Introduction
Additive manufacturing (AM), commonly known as 3D printing, is a transformative approach to industrial production. Laser powder bed fusion (LPBF) is the most common metal AM process. It has several advantages in comparison to conventional manufacturing technologies like milling and casting. LPBF offers design flexibility, quick fabrication of prototypes directly from digital models, create less material waste and demonstrate onsite manufacturing capability. The process of using LPBF to create a 3D object starts with spreading a layer of metal particles on the build plate. Then the laser source moving at a certain power and velocity are guided by the laser scanners according to the information in the 3D CAD to melt the regions relating to the design, which then solidifies to form a single layer. The powder spreading, melting and solidifying processes will occur repeatedly until the final object is formed. LPBF has attracted applications in key economic sectors such as in medicine, automotive, aerospace, construction, energy and prototyping [1, 2]. However, when compared to other form of manufacturing such as machining, casting and rolling, AM currently has the lowest market share, which is as a result of poor repeatability, defect formations, and low productivity [1, 3]. Issues with poor repeatability can be attributed to various process parameters such as scanning speed, laser power, beam radius, gas flow, layer thickness, and hatch spacing involved in operating LPBF. Slight errors in these parameters can propagate, resulting in variability in part properties. Additionally, the multi-physics phenomena such as heat transfer, solidification, melting, vaporization and cooling that occur in AM further makes it a sensitive process [4, 5, 6].
@@
Defects in LPBF could result from the process parameters that create the meltpool, and the common types are pore [7, 8, 9, 10, 11, 3, 12] and crack [13, 14, 15]. Pore defect can be divided into gas, lack-of-fusion (LOF) and keyhole-induced pore, and are differentiated by their natures and formation mechanism [16]. Spatter can lead to defects, especially to lack of fusion [13, 14, 15, 17]. Desire to implement multi-lasers to increase the productivity has been inhibited by spatter generation; an offshoot from the meltpool [18]. Their formation involves complicated interactions among high-speed vapor, vapor-induced powder motions and melt pool dynamics. Spatter can be the source of roughness and porosity on a part, and its detrimental effect could be reduced by manipulating the inert gas flow within the machine. However, reducing the quantity of spatter travelling during LPBF process requires optimum inert gas flow rate. High flow rate could lead to spatter generation from the powder bed, such that powder could be lifted by the inert gas, travel around the process chamber and land on a random location based on its trajectory, causing non-uniformity in the arrangement of the powder bed. It is challenging to establish optimum inert gas flow rate for spatter removal, also considering that the relationship will have to be established for varieties of power/velocity combinations which is not only time consuming but also expensive [5, 19, 14, 20]. Providing solution to spatter issues will require a holistic understanding of the physics of laser-material interactions, from melt pool creation to how the material properties can influence the intrinsic nature of the meltpool, leading to the instabilities that cause defect generation.
@@
The two general approaches for investigating different phenomena in AM are advanced observational technologies and numerical simulation methods. Few experimental studies have been conducted on spatter, Barrett et al. [21] used high speed sterovision camera to capture spatter, then used computer vision to collect their properties for statistical analysis. Previous work by Chen et al. [13] revealed five categories of spatter that can be formed during LPBF: solid, metallic jet, powder agglomeration, entrainment melting spatter, and induced spatter. Additionally, they found that the size, speed and direction of the spatters vary based on the process condition. Nakamura et al. [22] also recorded the molten pool using in-situ X-ray monitoring with the support of tungsten carbide as tracers in high power laser welding of titanium alloy. They stressed that the upward melt flow around the keyhole surface drove liquid column formation, which could cause spatter generation. Qu et al. [23] report the elimination of large spatters through controlling laser-powder bed interaction instabilities by using nanoparticles. The elimination of large spatters results in 3D printing of defect lean samples with good consistency and enhanced properties. Although advanced monitoring methods have been used to investigate the formation mechanism of spatters through observation of the physical process, the quantitative relationship between spatter formation and the physical process is still difficult to establish due to the experimental limitations of resolution and dimensions. A trial-and-error experimental procedure for determining appropriate processing parameters is extremely time-consuming and costly, and it remains challenging to completely monitor the complex physical phenomena occurring in LPBF at microsecond and micrometer scales.
@@
Numerical simulation can provide insights into regions where measurements are either difficult to collect or impossible. Additionally, it can offer three-dimensional, transient views of keyhole and molten pool behaviors that cannot be directly observed experimentally. Deng et al. [24] established a model of remote laser spiral spot welding to study the spatter formation. They found that the keyhole opening area serving as a main out-gassing channel and instantaneous vapor pressure impacting on the interface which are highly relevant with the spatter formation. Currently, many commercial simulation packages are available for AM, such as Ansys, Comsol multiphysics, and StarCCM+. However, these software lacks all the desired physical phenomenal that occur during AM. The available AM specific commercial software like FLOW-3D still require adjustments of various user numerical parameters, which sometimes are not accessible to the end-users. An open-source software like OpenFOAM (Open Field Operation and Manipulation) which is based on C++, is flexible in implementing additional features to address novel problems and processes occurring during AM process [25, 26].
@@
In recent years, machine learning (ML) have gained increasing attention in AM due to its unprecedented performance in sub-category tasks, such as supervised, un-supervised and reinforcement learning [2, 27, 28, 29, 30]. In addition, the high computational cost associated with high-fidelity simulations has motivated their synergy with ML tools, especially where a large scale of simulation experiments is necessary, such as in understanding defect formation processes, part design optimization, and the discovery of metamaterials and process monitoring. [31, 32, 33, 34, 35, 36, 37, 38]. Hemmasian et al. [39] used FLOW-3D simulation software to create datasets for LPBF processes and trained them using a convolutional neural network to predict the three-dimensional temperature field solely by taking the process parameters and the time step as input. DebRoy et al. [40] combined physics-informed machine learning, mechanistic modeling, and experimental data to identify several important variables that reveal the physics behind the defect formation in AM, by considering an example of balling defects in LPBF using peer-reviewed experimental data available in the literature. It is therefore possible to extend AM modelling and simulation to a large scale and to cover a wide range of parameters. Ren et al. [11] used x-rays to track the formation of pores in LPBF while also making observations with a thermal imaging system. This setup allowed the authors to develop a high-accuracy method for detecting pore formation from thermal signature with the help of a machine learning method.
The potential for the synergy of ML with physics-based simulation to understand spatter formation remains unexploited, which could have been due to limited studies and understanding of spatter as compared to other form of defects. Spatter generation is influenced by complex and often nonlinear interactions between multiple process variables. ML models are adept at capturing these complex relationships in high-dimensional data. ML can provide predictive insights that are not easily discernible through traditional physics-based modeling. It can predict when and under what conditions spatter is likely to occur. Lastly, ML can help in optimizing the process parameters not just to avoid spatter but also to achieve the best possible outcomes in terms of speed, cost, and material properties. By employing ML models, a deeper understanding of spatter can be achieved leading to enhanced control, improved process efficiency, and superior material characteristics in manufacturing and other material processing fields. In this study, we develop high fidelity multi-physics model based on OpenFOAM simulation to study spatter ejection from the meltpool in LPBF. Hot spatter is known to cause detrimental effects on the build component and act as a source of porosity; it is also known to originate from the melt pool. A tracking algorithm is developed to detect and study the behavior of ejected particles in three dimensions. We unveil the relationship between the spatter ejection and the meltpool via statistical analysis. Upon coupling ML model with multiphysics models, for the first time we reveal importance of different parameters to the spatter ejection and formulate spatter formation mechanism. Furthermore, our ML model was applied and tested against datasets procured from FLOW-3D, a prevalent yet computationally economical software extensively employed in additive manufacturing. However, FLOW-3D possesses inherent limitations, particularly an inability to correctly simulate ejected spatter motion directly due to restrictive assumptions in its gas phase model. By augmenting FLOW-3D with the ML model, we facilitate rapid and accessible quantification of generated spatter relative to the expected volume. This breakthrough led to the formulation of a comprehensive process map with insight to the spatter formation mechanism in additive manufacturing, which could serve as an excellent visualization tool to identify the process windows and optimize process parameters to print parts with low porosity and good surface finish.
@@
Figure 1: Spatter ejection and meltpool parameters curated from the OpenFOAM simulation are used as an input into a machine learning model for classification tasks. Trained machine learning models were used to test the dataset from the FLOW-3D, a commercial software, that lacks capability to simulation spatter formation, but less computationally expensive that it is 100 times faster in simulating meltpool behavior when compared to OpenFOAM. Based on the spatter prediction on FLOW-3D datasets, the process map for spatter in LPBF was generated to understand the root cause of defects as the process parameter varies.
@@ 2.1 Simulation of LPBF via OpenFOAM
Using OpenFOAMv2012, a CFD model was developed to understand the complex physical phenomena occurring during the LPBF process. A newly introduced icoReactingMultiphaseInterFoam (IRMIF) solver was used [1, 41, 25]. The IRMIF solver is based on the volume of fluid (VOF) method, where each phase is immiscible and a clear boundary between each phase is calculated. IRMIF solver addresses fluid mechanics including turbulent flow, heat transfer, laser beam sources with arbitrary beam shapes and phase transitions such as melting, solidification and evaporation. However, the solver did not account for the Marangoni effect due to the absence of the tangential component of the surface tension. The presence of the tangential component could influence the shape of the melt pool, leading to denudation, spatter and pore formations. Tangential component of the surface tension was added into the solver by Wirth et al. [18] following Brackbill equation [42]. as derived in equation (1)
$F_{s}=[\sigma kn+(\nabla\sigma-n(n\cdot\nabla\sigma))]\cdot\frac{2\cdot\rho(x)}{\rho_{1}+\rho_{2}}\cdot\frac{\left|\nabla_{\rho}(X)\right|}{\rho_{2}-\rho_{1}}$ (1)
Some assumptions were made such as: (1) the flow in the meltpool is Newtonian and turbulent using Reynolds-Averaged Simulation (RAS) (2) The thermophysical properties are of the function of temperature (3) plasma effect is ignored. Using the VOF method, the dynamic geometry of the free surface (gas/ metal interface) was captured. The Volume fraction satisfies equation (2)
$\frac{\delta\alpha_{L}}{\delta t}=-\vec{\nabla}\cdot\left(\alpha_{L}\vec{u}\right)-\vec{\nabla}\cdot\vec{u}\alpha_{L}\left(1-\alpha_{L}\right)$ (2)
@@
Where $\vec{u}$ is the flow velocity, $t$ is the time, and $\alpha_{L}$ is the liquid phase fraction. The finite volume method was used to compute the fluid flow, based on continuity and transient Navier stokes equation as described in equation (3) and (4) respectively.
$\frac{\partial\rho}{\partial t}+\nabla\cdot(\bar{\rho}\vec{U})=0$ (3)
$\frac{\partial(\rho\underline{u})}{\partial t}+\nabla(\rho\underline{u}\otimes\underline{u})=\nabla\left\{-p\cdot\underline{I}+\mu\left[\nabla\underline{u}+(\nabla\underline{u})^{T}\right]\right\}+\underline{F}$ (4)
where $\rho$ is the density, $\underline{u}$ is the fluid flow velocity vector, $t$ is the time, $p$ is the pressure, $I$ is the identity tensor, $\mu$ is the dynamic viscosity, and $\underline{F}$ is the volume force vector. The energy equation in equation (5) was used to compute the temperature field.
$\rho\frac{\delta E}{\delta t}=-\rho\vec{\nabla}\cdot\vec{u}E+\vec{\nabla}\cdot(\lambda\vec{\nabla}T+\overline{\bar{\tau}}\cdot\vec{u})\pm S_{h}$ (5)
where $E$ is the mixed energy, $\lambda$ is the thermal conductivity. IRMIF uses the Lee mass transfer model in equation (6) to calculate the solidification (or also melting) process, better described in [43] and the original formulation of Lee [44]

### v2

@@ Abstract
Laser powder bed fusion (LPBF) has shown promise for a wide range of applications due to its ability to fabricate freeform geometries and generate a controlled microstructure. However, components generated by LPBF still possess sub-optimal mechanical properties due to the defects that are created during laser-material interactions. In this work, we investigate the mechanism of spatter formation, using a high-fidelity modelling tool that was built using an open source platform know as OpenFOAM to simulate the multi-physics phenomena in LPBF. The model has the capability to capture the 3D resolution of the melt pool and the spatter behavior. Upon validating the OpenFOAM results with both experimental data and FLOW-3D, a commercial package, we found that the melt pool dimensions (length and width) were closely aligned across various conditions (150 W, 300 W, 450 W) at 1 m/s. The greatest deviation was a 7.8 % decrease in depth at 450 W from OpenFOAM to FLOW-3D, with other deviations falling under 4 %. To understand spatter behavior and formation, we reveal its properties at ejection and evaluate its variation from the melt pool, the source where it is formed. The dataset of the spatter and the melt pool collected consist of 50 % spatter and 50 % melt pool samples, with features that include position components, velocity components, velocity magnitude, temperature, density and pressure. The relationship between the spatter and the melt pool were evaluated via correlation analysis and machine learning (ML) algorithms for classification tasks. Upon screening different ML algorithms on the dataset, a high accuracy of 95 % and above was observed. Furthermore, the ML model was tested on a different dataset generated using FLOW-3D, a commercial, less computationally expensive tool, but lacks the capability to simulate realistic spatter formation. Our ML model trained on an OpenFOAM dataset was able to classify the region of the melt pool in the FLOW-3D datasets as spatter initiation location. Upon screening 281 FLOW-3D simulation experiments with varying power-velocity combinations, a process map based on spatter volume was established. The process map will serve as a guide towards understanding how the process parameters influence part properties, identifying an AM process window with minimal defects and ultimately enhance repeatability in part production.
@@ 1 Introduction
Additive manufacturing (AM), often referred to as 3D printing, is a transformative approach to industrial production. Laser powder bed fusion (LPBF) is the most common metal AM process. It has several advantages in comparison to conventional manufacturing technologies like milling and casting. LPBF offers design flexibility, quick fabrication of prototypes directly from digital models, reduced material waste, and demonstrates onsite manufacturing capability. The process of using LPBF to create a 3D object starts with spreading a layer of metal particles on the build plate. Then the laser source moving at a certain power and velocity are guided by the laser scanners according to the information in the 3D CAD to melt the regions relating to the design, which then solidifies to form a single layer. The powder spreading, melting and solidifying processes will occur repeatedly until the final object is formed. LPBF has attracted applications in key economic sectors such as medicine, automotive, aerospace, construction, energy, and prototyping [1, 2]. However, compared to other forms of manufacturing such as machining, casting, and rolling, AM currently has the lowest market share, which is due to poor repeatability, defect formation, and low productivity [1, 3]. Issues with poor repeatability can be attributed to various process parameters such as scanning speed, laser power, beam radius, gas flow path, layer thickness, and hatch spacing involved in operating LPBF. Slight changes in these parameters can propagate, resulting in variability in the properties of parts. Furthermore, the multi-physics phenomena such as heat transfer, solidification, melting, vaporization, and cooling that occur in AM further make it a sensitive process [4, 5, 6].
@@
Defects in LPBF could result from the process parameters that generate the melt pool, and the common types are pore [7, 8, 9, 10, 11, 3, 12] and crack [13, 14, 15]. Taherkhani et al. [16] used the in-situ photoanode sensor to analyze the light signal emitted from the melt pool to quantify the defect and found it to correlate with the process condition. Pores are classified into gas, lack of fusion (LOF) and keyhole induced as a result of their formation from gas entrapment, incomplete melting, or improper print settings, respectively [17]. Rollett et al. [18] used ultrahigh-speed synchrotron X-ray imaging to quantify vapor depressions during melting and identified a clear laser power density threshold for the transition from conduction to keyhole mode. They established quantitative relationships between laser power density, keyhole depth, and front wall angle, providing insight into predicting and controlling keyhole behavior. Spatter ejected from the melt pool is known to cause defects such as surface roughness, porosity, and inclusions [13, 14, 15, 19]. Snow et al. [20] demonstrated that spatter particles play a direct role in causing stochastic lack-of-fusion defects in LPBF components. Using spatial statistics, they revealed that the locations of internal flaws are intrinsically connected to the distribution of spatter particles. Toyserkani et al. [21] found that spatter increased surface roughness from 14.4 $\mu m$ to 28 $\mu m$, recommending avoiding spatter-rich regions for high-quality parts. The desire to implement multi-lasers to increase printing productivity has been inhibited by spatter generation [22]. Spatter formation results from complex interactions between high-speed vapor, vapor-induced powder movement, and melt pool dynamics. The detrimental effect of spatter could be reduced by manipulating the inert gas flow within the machine. However, reducing the amount of spatter traveling during the LPBF process requires an optimum inert gas flow rate. The high flow rate could lead to spatter generation from the powder bed, such that powder could be lifted by the inert gas, travel around the process chamber and land on a random location based on its trajectory, causing non-uniformity in the arrangement of the powder bed [23]. It is challenging to establish optimum inert gas flow rate for spatter removal, also considering that the relationship will have to be established for varieties of power/velocity combinations which is not only time consuming but also expensive [5, 24, 14, 25]. Providing solution to spatter issues will require a holistic understanding of the physics of laser-material interactions, from melt pool creation to how the material properties can influence the intrinsic nature of the melt pool, leading to the instabilities that cause defect generation across wide process condition space.
@@
The two general approaches to investigating different phenomena in AM are advanced observational technologies and numerical simulation methods. Few experimental studies have been conducted on spatter, Barrett et al. [26] used a high-speed sterovision camera to capture the spatter, then used computer vision to collect their properties for statistical analysis. Previous work by Chen et al. [13] revealed five categories of spatter that can be formed during LPBF: solid, metallic jet, powder agglomeration, entrainment melting spatter, and induced spatter. Additionally, they found that the size, speed, and direction of the spatters vary based on the process condition. Nakamura et al. [27] investigated the behavior of the molten pool during high-power laser welding of titanium alloy using in-situ X-ray monitoring. They highlighted that the upward melt flow around the keyhole surface contributed to the formation of a liquid column, potentially causing spatter generation. Qu et al. [28] reported that coating Al6061 with TiC nanoparticles effectively eliminated large spatter generation, demonstrating the ability of the nanoparticles to control instabilities during laser-powder bed interactions. This approach resulted in parts with less defects, better consistency, and improved properties. Although advanced monitoring techniques have been used to investigate the mechanism of spatter formation via the physical process, establishing a quantitative relationship between spatter formation and the physical process remains challenging due to experimental limitations in resolution and scale. Determining suitable processing parameters using a trial-and-error approach is time-consuming and expensive. Additionally, it is challenging to fully monitor the complex physical phenomena in the LPBF at microsecond and micrometer scales.
@@
Numerical simulation can offer valuable insight into areas where measurements are challenging or impossible to obtain. It also provides three-dimensional, time-dependent views of keyhole and molten pool behaviors that cannot be directly observed through experiments. Deng et al. [29] developed a model of remote laser spiral spot welding to investigate spatter formation. Their findings revealed that the opening area of the keyhole acts as a primary out-gassing channel, and the instantaneous vapor pressure on the interface is closely related to spatter formation. Many commercial simulation packages such as Ansys, Comsol multiphysics, and StarCCM+ are used for investigating AM. The available AM specific commercial software like FLOW-3D still require adjustments of various numerical parameters, which sometimes are not accessible to the end-users. An open-source software like OpenFOAM (Open Field Operation and Manipulation), is flexible in implementing additional features to address novel problems and processes occurring during AM process [30, 31]. However, modeling AM process by taking into account the multiphysics involved in the melt pool dynamics and spatter generation is computationally expensive.
@@
The high computational cost associated with high-fidelity simulations has motivated their synergy with ML tools, especially where a large scale of simulation experiments is necessary, such as in understanding defect formation processes, part design optimization, and the discovery of metamaterials and process monitoring [2, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42]. Hemmasian et al. [43] utilized FLOW-3D simulation software to generate data sets for LPBF processes and trained a convolutional neural network to predict the temperature field using only the process conditions as input to the model. DebRoy et al. [44] integrated physics-informed machine learning, mechanistic modeling, and experimental data to identify key variables that shed light on the physics behind defect formation in AM. Using balling defects in LPBF as an example, they utilized peer-reviewed experimental data from the literature. This approach makes it possible to extend AM modeling and simulation to a larger scale and encompass a wider range of parameters. Ren et al. [11] employed X-rays to monitor pore formation in LPBF while simultaneously using a thermal imaging system for observations. This setup enabled the authors to develop a highly accurate method for detecting pore formation from thermal signatures, utilizing a machine learning approach.
The potential for synergy of ML with physics-based simulation to understand spatter formation remains unexploited, which could have been due to limited studies and understanding of spatter as compared to other forms of defects. Spatter generation is influenced by complex and often non-linear interactions between multiple process variables. ML models are adept at capturing these complex relationships in high-dimensional data [41, 35]. ML can provide predictive insights that are not easily discernible through traditional physics-based modeling. Lastly, ML can help optimize process parameters not only to avoid spatter but also to achieve the best possible results in terms of speed, cost, and material properties [15]. Using ML models, a deeper understanding of the spatter can be achieved, leading to improved control, improved process efficiency, and superior material characteristics in manufacturing and other material processing fields.
In this study, we develop a high fidelity multi-physics model based on OpenFOAM simulation to study the spatter ejection in LPBF. Hot spatter is known to cause detrimental effects on the build component, act as a source of porosity, and roughness as it can create local melting upon landing on the solid part as it is known to originate from the melt pool. Compared to powder spatter, the formation of hot spatter and its role in inducing macroscopic defects have a more significant impact on the quality of the LPBF process [45]. Therefore, our work aims at developing a methodology specifically for analyzing the properties that differ between the ejected fluid and the melt pool, as detecting fluctuations in the melt pool is critical for ensuring satisfactory quality control during part fabrication. Having established these studies on the behavior of hot spatter ejection, not only will misclassification of the different types of spatter be prevented, their effect will be captured and deconvoluted in the future investigation. We develop a tracking algorithm specifically for the LPBF dataset generated with OpenFOAM, which differentiates between spatter and the melt pool, links data across consecutive time steps, and extracting properties at each time step, such as temperature, velocity, location, and density in 3D. We unveil the relationship between the spatter ejection and the melt pool via statistical analysis, whereby spatter is detected at a short time scale (5 µs), paving the way to quantifying the fluctuation in the melt pool for good quality control. Upon coupling our ML model with multiphysics models, for the first time we reveal importance of different parameters to the spatter ejection and formulate a spatter formation mechanism. In addition to using ML models for classification tasks and to establish trends in the feature importance, we complement it with an explainable AI technique which provide insights into the range of the feature values that positively contribute spatter predictions. Furthermore, our ML model was applied and tested against datasets generated from FLOW-3D, a prevalent yet computationally economical software extensively employed in AM. However, FLOW-3D implements restrictive assumptions that limit spatter simulation to improve computational cost. Specifically, the gas dynamics within the keyhole are not directly simulated, and are instead approximated by estimates of the recoil pressure and mass transfer. Additionally, FLOW-3D lacks the tangential surface tension component [46, 47]. In contrast, our OpenFOAM model keyhole gas dynamics and implements tangential surface tension following the method by Wirth et al. [1]. By leveraging the respective strengths of OpenFOAM and FLOW-3D simulations, augmented with machine learning (ML), a reduced-order methodology was developed to accelerate spatter process map generation in LPBF. This approach facilitates rapid and accessible quantification of generated spatter relative to the expected volume. This enables the creation of a comprehensive process map, providing insight into the spatter formation mechanism in additive manufacturing. This map can be used as a visualization tool to identify optimal process windows and parameters, enabling the printing of parts with low porosity and high-quality surface finish.
@@
Figure 1: Spatter ejection and melt pool parameters from OpenFOAM simulations were used as input for machine learning classification tasks. The trained models were tested on FLOW-3D datasets, a commercial software that lacks realistic spatter simulation capability to generate a process map. A 5 $\mu m$ LPBF simulation in FLOW-3D and OpenFOAM requires 4 hours and 72 hours respectively on the same machine.
@@ 2.1 Simulation of LPBF via OpenFOAM
A CFD model was developed using OpenFOAMv2012 to analyze the complex physical phenomena that occur during the LPBF process. A newly introduced icoReactingMultiphaseInterFoam (IRMIF) solver [1, 48, 30] was used. The IRMIF solver is based on the volume-of-fluid (VOF) method, where each phase is immiscible, and a clear boundary between each phase is calculated. The IRMIF solver handles fluid mechanics, including turbulent flow, laser beam sources with arbitrary beam shapes, heat transfer, and phase transitions such as solidification, melting, and evaporation. However, the solver did not account for the Marangoni effect due to the absence of the tangential component of the surface tension. The presence of the tangential component could influence the shape of the melt pool, leading to denudation, spatter, and pore formations. The tangential component of the surface tension is added to the solver following the modification by Wirth et al. [22] to incorporate the Brackbill equation [49] which enables the implementation of surface tension forces in the VOF method as shown in equation (1)
$F_{s}=\left[\sigma\kappa\vec{n}+\left(\nabla\sigma-\vec{n}(\vec{n}\cdot\nabla\sigma)\right)\right]\cdot\frac{2\cdot\rho(x)}{\rho_{1}+\rho_{2}}\cdot\frac{\left|\nabla\rho(x)\right|}{\rho_{2}-\rho_{1}}$ (1)
The parameters in Equation (1) include the surface tension $\sigma$, the interface curvature $\kappa$, the normal vector of the interface $\vec{n}$, and the densities $\rho_{1}$ and $\rho_{2}$ of the two phases at the interface. In brackets of Equation (1), the first term corresponds to the normal component of the surface tension force, while the second term corresponds to the tangential component. The factor $\frac{2\cdot\rho(x)}{\rho_{1}+\rho_{2}}$ acts as a weight function that redistributes the surface tension force smeared to the denser phase. The term $\frac{\left|\nabla\rho(x)\right|}{\rho_{2}-\rho_{1}}$ determines the position of the interface where it is non-zero.
Using the VOF method, the dynamic geometry of the free surface (gas/ metal interface) was captured. The volume fraction satisfies equation (2). In OpenFOAM, an additional convective term is introduced to the VOF, adapting it to a two-fluid Eulerian system [50], which depends on the compression velocity $\vec{n}$. This velocity is active only in the interface cells due to the term $\alpha_{1}(1-\alpha_{1}$). This artificial velocity helps compress the interface by reducing the numerical diffusion of the volume fraction $\alpha_{1}$ while preserving its boundedness [51]. Since only the VOF is solved by the solver used in this study, the forces at play drive the fluid flow, with the velocity determined through the momentum balance to account for ejection from the melt pool. Since the force acting on the particles is intrinsic to the fluid flow, as it is a viscous fluid rather than individual particles, solving the velocity field captures the dynamics of spatter ejection.
[…]
In the situation of the mixture of vapor and nitrogen, the model applies a weighting average based on the composition of each component in the mixture in equation (3-7) to the value of its physical properties, such as density $\rho$, viscosity $\mu$, thermal conductivity $k$ and specific heat capacity $C$ as listed in Table 1.
$\alpha_{1}+\alpha_{2}=1$ (3)
$\rho=\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2}\$ (4)
$\mu=\alpha_{1}\mu_{1}+\alpha_{2}\mu_{2}\$ (5)
$k=\alpha_{1}k_{1}+\alpha_{2}k_{2}\$ (6)
$C=\alpha_{1}\frac{\rho_{1}}{\rho}C_{1}+\alpha_{2}\frac{\rho_{2}}{\rho}C_{2}$ (7)
@@
The value of either $\alpha_{1}$ or $\alpha_{2}$ ranges from 0 to 1, where $\alpha_{1}$ and $\alpha_{2}$ are the volume fractions of the phase of the element, $\vec{u}$ is the flow velocity, $t$ is the time. The finite volume method was used to compute the fluid flow, based on continuity and transient Navier-Stokes equation, as described in equations (8) and (9) respectively.
$\frac{\partial\rho}{\partial t}+\nabla\cdot(\rho\vec{u})=0$ (8)
$\frac{\partial(\rho\vec{u})}{\partial t}+\nabla\cdot(\rho\vec{u}\otimes\vec{u})=\nabla\cdot\left\{-p\cdot\vec{I}+\mu\left[\nabla\vec{u}+(\nabla\vec{u})^{T}\right]\right\}+\vec{F}$ (9)
where $\rho$ is the density, $\vec{u}$ is the fluid flow velocity vector, $t$ is the time, $p$ is the pressure, $I$ is the identity tensor, $\mu$ is the dynamic viscosity and $\vec{F}$ is the volume force vector. Rapid heating to high temperatures causes the molten material to evaporate, generating a recoil pressure within the melt pool. This recoil pressure, produced by the vapor jet, drives the melt surface downward, creating a vapor depression. The recoil pressure is modeled using the Clausius–Clapeyron equation to calculate the saturated vapor pressure, adjusted by a coefficient that accounts for the backward flux of the evaporated material, as shown in equation 10 [52, 53, 54].
$\vec{P}_{r}=0.54P_{0}\exp\left[\frac{L_{v}M(T-T_{v})}{RTT_{v}}\right]\vec{n}\left|\nabla\alpha_{1}\right|\frac{2{\rho}}{\rho_{\text{metal}}+\rho_{\text{gas}}}$ (10)
where $P_{0}$ is the ambient pressure, $L_{v}$ is the latent heat of evaporation, $T_{v}$ is the evaporation temperature, $M$ is the molar mass, and $R$ is the universal gas constant. The energy equation in equation (11) was used to compute the temperature field.
$\rho\frac{\partial E}{\partial t}=-\rho\vec{\nabla}\cdot\vec{u}E+\vec{\nabla}\cdot(k\vec{\nabla}T+\overline{\bar{\tau}}\cdot\vec{u})+q_{v}+q_{\text{rad }}+q_{\text{laser }}$ (11)
where $E$ is the mixed energy, $k$ is the thermal conductivity, $T$ is the temperature, $\overline{\bar{\tau}}$ is the stress tensor. $q_{\text{v}}$ is the heat loss due to vaporization (Equation 12), $q_{\text{rad}}$ is the heat loss due to radiation (Equation 13), $q_{\text{laser}}$ is the heat input from the laser beam (equation 14).
$q_{v}=-0.82\frac{l_{v}M}{\sqrt{2\pi MRT}}P_{0}\exp\left[\frac{l_{v}M\left(T-T_{v}\right)}{RTT_{v}}\right]\left|\nabla\alpha_{1}\right|\frac{2\rho C_{p}}{\rho_{\text{metal }}\rho C_{\text{pmetal }}+\rho_{\text{gas }}C_{\text{pgas }}}$ (12)
$q_{\text{rad }}=-\sigma_{s}\varepsilon\left(T^{4}-T_{\text{ref }}^{4}\right)\left|\nabla\alpha_{1}\right|\frac{2\rho C_{p}}{\rho_{\text{metal }}\rho C_{\text{pmetal }}+\rho_{\text{gas }}C_{\text{pgas }}}$ (13)
Where Where $C_{p}$ is the specific heat capacity, $L_{m}$ is the latent heat of melting, $\sigma_{s}$, $\varepsilon$, $T_{\text{ref }}$ are the Stefan-Boltzmann constant, emissivity and ambient temperature respectively. The IRMIF solver uses the discrete transfer radiation model, which takes the laser beam into account, passing through VOF phases [55]. The laser energy was applied in the form of a body heat source and goes into the depth of the material. A Gaussian intensity distribution $(D_{4\sigma}=100\,\mu\text{m})$ is applied to the laser beam.
$q_{\text{laser }}=\frac{f_{\text{absorb }}I}{\Delta y}\quad\text{ where }I=\frac{2\xi P}{\Pi r^{2}}\exp\left[\frac{-2\left(x+v_{0}t-x_{0}\right)^{2}+\left(z-z_{0}\right)^{2}}{r^{2}}\right]$ (14)
Where $I$ is the beam power density, $f_{\text{absorb }}$ is the percentage of laser energy occupied by the element, $P$ is the laser power, $\xi$ is the absorption coefficient, $r$ is the beam radius, ($x_{0}$, $z_{0}$) describes the starting position of the laser beam, $v_{0}$ denotes the scanning velocity and $\Delta y$ represents the size of the numerical grid. IRMIF uses the Lee mass transfer model in equation (15) to calculate the solidification (or also melting) process, better described in [56] and the original formulation of Lee [57].